m/5 + 8 = -7
I am tutoring a student working on his GED and need help with this problem.
Thank you in advance for any help you can provide.
Cherryl

Answer by checkley75(3666)

M/5+8=-7 SUBTRACT 8 FROM BOTH SIDES OF THIS EQUATION
M/5+8-8=-7-8
M/5=-15 NOW CROSS MULTIPLY
M=-75 ANSWER.
PROOF
-75/5+8=-7
-15+8=-7
-7=-7